a rectangle is 3 times as long as it is wide. the perimeter is 60 cm. find the dimensions of the rectangle. round to the nearest tenth if necessary
Answer by Jeannie1012(8)

The formula for finding out the perimeter of a rectangle is 2L+2w= perimeter
In this problem we are told that the perimiter is = to 60 and that the length is 3 times as long as the width. So L =2(3w)
your equation is 2(3w)+2w=60
use distribution to solve the 2(3w) and get 6w+2w=60
combine like terms 8w=60
divide both sides by 8 to get w alone 8w divided by 8 is w and 60 divided by 8 is 7.5 so w= 7.5
now that you know w is = to 7.5 and you know that the lenth is 3 times long you mulitply 7.5 by three and get 22.5
w=7.5
L=22.5
to check plug the numbers back in 2(7.5) + 2(22.5) = 60? 15+ 45 = 60 since it checks its correct.